The graph of y = n ( 1−x^2 ) has an area equal to 51, find the value of n.
Answer by rothauserc(4718)

first find the zeros for 1 - x^2
(1+x)(1-x) = 0
zeros are -1, 1
integrate 1 - x^2
integral = x - x^3/3
evaluate the integral for x = -1 and 1
(1 - 1/3) - (-1 + 1/3) = 2/3 + 2/3 = 4/3
we are given
51 = n * 4/3
4n = 153
n = 38.25